OpenStudy (anonymous):

Help me with this problem please:

OpenStudy (anonymous):

\[\sqrt{5x}\]

hartnn (hartnn):

that is the most simplified form all u can do is: \(\sqrt{ab}=\sqrt{a}\sqrt b\)

OpenStudy (anonymous):

Yes, he is right, this is already in the simplest form...

OpenStudy (anonymous):

tHANKS gUYS!